OpenStudy (anonymous):

trigonometry 3) Simplify. a)sec x/tan x b)(sec - 1) (sec x + 1)

OpenStudy (anonymous):

The first one just write sec x as 1/cosx and write tan x as sinx/cosx So now just divide 1/ cos(x) by sin(x)/ cos(x) So invert and multiply. 1/ cos(x) * cos(x)/sin(x) = 1/sin(x) 1/sin(x) = csc(x) For b, note you have a difference of squares. So (sec(x) -1)(sec(x)+1) = sec^2(x) -1 Write sec^2(x) as 1/cos^2(x) So now you have this: 1/cos^2(x) - 1 Find a common denominator and rewrite equivalent fractions. 1/cos^2(x) -cos^2(x)/cos^2(x) Note that 1-cos^2(x) is the Pythagorean identity and equals sin^2(x) So we now have this: sin^2(x)/cos^2(x) = tan^2(x)